Given a sample size of 14, what are the critical values for a 98% confidence level?
Answer #1

solution Given that n14 degrees at facedom h-1 =14-1 13 At 98.% Contidence level th tis azI-98 =1-0.9& O O2 2/2- 0.01 ty2,dft